OpenStudy (anonymous):

ae 7x-4=10 and x=3 eqivalent equations? explain

OpenStudy (anonymous):

well that looks wrong wouldn't x=2? are you sure you mean x=2 instead of x=3?

OpenStudy (anonymous):

anyhoo if i remember right equivilent equations is a linear algebra concept. used to make problem easier to solve You can do things to equations like this x=y x+c=y+c add TBS x*c=y*c multiply BS and so on and so on.